OpenStudy (anonymous):

A right triangle has legs that are 15 centimeter long and 7 centimeters long. What is the length of the hypotenuse? i know this is pythagoream theorem. And the 15 is under a radical sign.

OpenStudy (accessdenied):

So, \( a^2 + b^2 = c^2 \), where \( a = \sqrt{15} \) and \( b = 7 \)?

OpenStudy (anonymous):

oh right. uhm. It's the legs.

OpenStudy (anonymous):

How would I solve the radical 15?

OpenStudy (accessdenied):

yeah, \(a\) and \(b\) would be the lengths of the legs and \(c\) would be the hypotenuse. Then we just solve for c by evaluating the left side and then taking its square root.

OpenStudy (accessdenied):

\( (\sqrt{15})^2 = 15 \)

OpenStudy (anonymous):

Ohhh. I see So the answer would be about seventeen? The choices are 5, 8, 17, and 22

OpenStudy (accessdenied):

\( (\sqrt{15})^2 + 7^2 = c^2\\ 15 + 49 = c^2\\ 64 = c^2 \)

OpenStudy (anonymous):

ohhh! so you don 't need to put 15 to the power of 2 i see. Thanks so much. :3

OpenStudy (accessdenied):

No problem. :) We just substitute the values we have into the Pythagorean theorem, and square roots will tend to disappear in the lengths because the lengths of the legs are always squared in that formula.

OpenStudy (anonymous):

Thank you. I'm such a blonde...even though I have black hair :P

OpenStudy (accessdenied):

Haha, I had a blonde moment (brown hair of course) today doing a sequence problem on a test. My mind went completely blank on how to do it until after the test, when my friend told me how he did it. :P